Problem 1. D= C-S/N (for S)

Problem 2. A=1/2h(B+b)(for b)

Answer by malakumar_kos@yahoo.com(315)


Problem 1. D= C-S/N (for S)
Problem 2. A=1/2h(B+b)(for b)
1) D = C-S/N
-S/N = D-C
-S = N(D-C)
S = -N(D-C) = N(C-D)
S = N(C-D)


2)A = 1/2 h(B+b)
2A = h(B+b)
2A/h = B+b
2A/h - B = b Therefore b = (2A/h)-B .